A particle is moving along a straight line such that its acceleration can be given as a=3-4t, where t is in second. At t=0, its velocity is v=2m/s. Distance travelled by particle in 3 s is

Grade:11

1 Answers

Arun
6 years ago
a = 3 -4t
Hence dv = a.dt
dv = (3-4t) dt
On integrating
v = 3t - 2t² +c
But at t= 0, v =2 hence
C = 2
Now
ds = v dt
On integrating
S = -(2/3)t³ +(3/2) t² +2t
Hence
Distance travelled in 3 sec
put t = 3
S = -18 +13.5 +6
S = 1.5 m